Grade 12th passGeneral Physics

a simple pendulum with bob of mass m and length is held in position at an angle o and then angle o2 with vertical when released from these positions speed with which it passes lowest position are v1 and v2 then v1/v2 is

harshita
8 Years agoGrade 12th pass

1 Answer

Babli kumari
8 Years ago
Height of pendulum raised,=x-xcosπ1=x(1-cosπ)Now the PE=mgh=mgx(1-cosπ1)Now the KE=1/2mv^21/2mv^2=mgx(1-cosπ1).......aSimilarly for second position1/2mv^2=mgx(1-cosπ2).......bNow from and bV1/v2=√1-cosπ1/√1-cosπ2